40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 18:35:09 2019, gautam, Update, LSC, Locking prep 
    Reply  Wed Oct 9 18:15:26 2019, gautam, Update, LSC, PRMI 3f locking PRMI_1f.png
       Reply  Wed Oct 9 22:02:58 2019, gautam, Update, LSC, REFL55 whitening issue 
          Reply  Thu Oct 10 01:12:56 2019, gautam, Update, LSC, Locking studies PRFPMIstudies.png
             Reply  Wed Oct 16 11:42:17 2019, gautam, Update, LSC, Poor separation of PRCL/MICH in 3f signals sensMat.pdf
                Reply  Thu Oct 17 11:19:28 2019, gautam, Update, LSC, Locking activity last night MICH2PRCLnulling.pdfconsolidatedSensingMatrices.pdf.zip
                   Reply  Tue Oct 22 00:52:27 2019, gautam, Update, LSC, Locking updates PRMI3f_ALS_Oct21sensMat.pdf
                      Reply  Tue Oct 29 01:24:45 2019, gautam, Update, LSC, More locking updates POP22_feature.pngPRMI3f_ALS_Oct21sensMat.pdf
                         Reply  Tue Oct 29 15:13:19 2019, gautam, Update, LSC, More locking updates PRMI_lockloss.png
                            Reply  Wed Oct 30 01:27:00 2019, gautam, Update, LSC, More locking updates POP_ASS.png
                               Reply  Wed Oct 30 11:53:41 2019, gautam, Update, LSC, MICH loop shape tuning MICH_ELP.pdfMICH_ELP_TFs.pdf
                               Reply  Wed Oct 30 17:08:40 2019, gautam, Update, LSC, POP22 investigation POP22anomaly.pdf
                                  Reply  Mon Nov 4 15:29:47 2019, gautam, Update, LSC, POP signal path POPchain.pdf
                                     Reply  Mon Nov 4 16:06:58 2019, gautam, Update, LSC, POP optical path IMG_8027.JPGlensRealignment.jpgPOPrealigned.pngPOPdither.pngPRMfixed.png
                                        Reply  Tue Nov 5 11:52:27 2019, gautam, Update, LSC, Locking notes PRMI3f_ALS_Nov4sensMat.pdf
                                           Reply  Wed Nov 6 02:08:48 2019, gautam, Update, LSC, Locking updates 
                                              Reply  Wed Nov 6 17:05:45 2019, gautam, Update, LSC, CARM calibration CARMscan.pdfCARMcalib.pdfscan.hdf5.zip
                                              Reply  Wed Nov 6 17:45:34 2019, gautam, Update, LSC, ~ 
                                                 Reply  Mon Nov 18 21:04:38 2019, gautam, Update, LSC, Locking - some ideas 
                                                    Reply  Wed Nov 20 01:07:18 2019, gautam, Update, LSC, Locking - progress PRFPMI.png
                                                       Reply  Wed Nov 20 12:14:17 2019, gautam, Update, LSC, Locking - progress TRX_TRY_comparison.pdf
                                                          Reply  Wed Nov 20 17:20:24 2019, Yehonathan, Update, LSC, QPD Investigation 
                                                             Reply  Wed Nov 20 17:52:00 2019, gautam, Update, LSC, QPD MEDM screen update newLookQPD.pngTRX_TRY_comparison.pdfIMG_8186.JPG
                                                                Reply  Wed Nov 20 21:29:28 2019, gautam, Update, LSC, PRG ~13 PRG13.pdf
                                                                   Reply  Fri Nov 22 00:54:14 2019, gautam, Update, LSC, locking notes ALSnoiseIncrease.pdf
                                                                      Reply  Mon Nov 25 19:11:22 2019, gautam, Update, LSC, ALS noise re-look ALSnoise.pdf
                                                                         Reply  Wed Nov 27 23:24:01 2019, gautam, Update, LSC, No shaking but no inspiration either 
                                                                            Reply  Sun Dec 1 13:38:48 2019, rana, Update, LSC, No shaking but no inspiration either 
                                                                               Reply  Tue Dec 3 00:03:57 2019, gautam, Update, LSC, Green ALS also shows elevated noise with high arm buildup ALSnoiseIncrease_greenBeat.pdf
                                                                                  Reply  Wed Dec 4 19:54:27 2019, gautam, Update, LSC, A look to the past DRFPMI_2016March.pdf
                                                                                     Reply  Thu Dec 5 01:54:39 2019, gautam, Update, LSC, Partial CM board path engaged AOengaged.pdf
                                                                                        Reply  Fri Dec 6 00:02:48 2019, gautam, Update, LSC, What is the correct way to set the 3f offsets? 
                                     Reply  Fri Nov 15 12:08:04 2019, gautam, Update, LSC, POPDC whitening board 
                         Reply  Tue Oct 29 17:40:48 2019, gautam, Update, LSC, More locking updates broadSweep.pdfzoomSweep.pdf
Message ID: 15040     Entry time: Wed Nov 20 17:52:00 2019     In reply to: 15039     Reply to this: 15041
Author: gautam 
Type: Update 
Category: LSC 
Subject: QPD MEDM screen update 

Koji and I had noticed that there was some discrepancy between the switchable gain stages of the EX and EY QPDs. Sadly, there was no indication that these switches even exist on the QPD MEDM screen. Yehonathan and I rectified this today. Both EX and EY Transmon QPDs now have some extra info (see Attachment #1). We physically verified the indicated quadrant mapping for the EX QPD (see previous elog in this thread for the details), and I edited the screen accordingly. EY QPD still has to be checked. Note also that there is an ND=0.4 + ND=0.2 filter and some kind of 1064nm light filter installed in series on the EX QPD. The ND filters have a net transmission T~25%.

After making the EX and EY QPDs have the same switchable gain settings (I also reset the trans normalization gains), the angular motion witnessed is much more consistent between the two now - see Attachment #2. The high-frequency noise of the sum channel is somewhat higher for EX than EY - maybe the ND filters are different on the two ends?

Note that there was an extra factor of 40 gain on the EX QPD relative to EY during the lock yesterday. So really, the signals were probably just getting saturated. Now that the gains are consistent between the ends, it'll be interesting to see how balanced the buildup of the two arms is. There still remains the problem that the MICH loop was too unstable, which probably led to to excess arm power fluctuations.

There is a mark on the QPD surface that is probably dirt (since we never have such high power transmitted through the ETM to damage the QPD). I'll try cleaning it up at the next opportunity.

Attachment 1: newLookQPD.png  75 kB  | Show | Hide all | Show all
Attachment 2: TRX_TRY_comparison.pdf  119 kB  | Hide | Hide all | Show all
TRX_TRY_comparison.pdf
Attachment 3: IMG_8186.JPG  822 kB  | Hide | Hide all | Show all
IMG_8186.JPG
ELOG V3.1.3-